OpenStudy (anonymous):

Please help?!! solve the equation √z-10=16. check your answer

OpenStudy (anonymous):

\[\sqrt{z}-10=16\] \[\sqrt{z}=16+10\] \[\sqrt{z}=26\] \[(\sqrt{z})^{2} =16^{2}\] \[z=256\]